Solve for x:

73x-1 =32x

1
Expert's answer
2021-11-08T17:34:28-0500

"7^{3x-1}=3^{2x}"

"(3x-1)ln(7)=2xln(3)"

"x=\frac {ln(7)}{3ln(7)-2ln(3)}"